INFRARED COMBINATION AND DIFFERENCE BANDS OF THE NO DIMER

KE-XIAN AU YONG, JENNA M. KING, \underline{A.R.W. McKELLAR}, and J.K.G. WATSON, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The mid-infrared (1500 - 3800 cm^{−1}) absorption spectrum of gaseous nitric oxide has been studied at low temperature (99 K) with a long absorption path (160 m) in order to observe weak combination, difference, and overtone bands of the NO dimer. Ten new bands were assigned with greater or lesser certainty. Combined with previous results, they lead to a set of 12 secure and 7 tentative vibrational term values for (NO)_{2}, which almost double our knowledge of NO dimer vibrational states. The strongest non-fundamental bands in this region, after the ν_{1} (symmetric N-O stretch) + ν_{5} (asymmetric N-O stretch) overtone, involve combinations of ν_{5} with ν_{3} (intermolecular stretch). A new value of 155.5 cm^{−1} was obtained for ν_{4}, the elusive infrared-inactive out-of-plane fundamental. Excitation of ν_{5} results in increased frequencies for the intermolecular modes ν_{2}, ν_{3}, and ν_{4}, consistent with the known strengthening of the intermolecular bond.